OpenStudy (anonymous):

Find the points of inflection; 2csc(3x/2) (0,2pi)..describe the concavity

myininaya (myininaya):

\[f(x)=2\frac{1}{\sin(3x/2)}\] \[f'(x)=2*\frac{0*\sin(3x/2)-3/2*\cos(3x/2)}{\sin^2(3x/2)}=-3\frac{\cos(3x/2)}{\sin^2(3x/2)}\] \[f''=3*\frac{-3/2*\sin(3x/2)\sin^2(3x/2)-2\sin(3x/2)*\cos(3x/2)*3/2*\cos(3x/2)}{\sin^4(3x/2)}\] \[f''=3*\frac{3}{2}*\sin(3x/2)*\frac{-\sin^2(3x/2)-\cos^2(3x/2)}{\sin^4(3x/2)}\] \[f''=\frac{-9}{2}\sin(3x/2)*\frac{1}{\sin^4(3x/2)}\] to find possible inflection points set \[\sin(3x/2)=0\]

myininaya (myininaya):

remember sin(0)=0 and sin(pi)=0 and sin(2pi)=0 but you said not include 0 and 2pi since you used () and not [] right? so anyways sin(pi)=0 so we set 3x/2=pi and solve for x

myininaya (myininaya):

x=2pi/3

myininaya (myininaya):

now remember this gives us that we have a possible inflection point at (2pi/3,f(2pi/3))

myininaya (myininaya):

we need to see if the concavity switches there if so then we indeed have an inflection point

myininaya (myininaya):

-------|---------- 2pi/3 choose a number btw 0 and 2pi/3 (3x/2=0 => x=0) choose a number btw 2pi/3 and 4pi/3 (3x/2=2pi=>x=4pi/3)

myininaya (myininaya):

plug those numbers into f'' a positive number=> concave up a negative number=> concave down

myininaya (myininaya):

oops so actually you have one more interval you need to choose a number btw 4pi/3 and 2pi actually you will not have any inflection points since f(2pi/3) and f(4pi/3) doesn't exist but still you use the above numbers to test for concavity